Assume that each atom in a copper wire contributes one free electron. Estimate the number of free electrons in a copper wire having a mass of 6.4 g (take the atomic weight of copper to be 64 g/mol).

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. 64 grams of copper have 1 mole 6.4 grams of copper have 0.1 mole 1 mole = No atoms 0.1 mole = (no × 0.1) atoms = 6 × 10^23 × 0.1 atoms = 6 × 10^22 atoms 1 atom contributes 1 electron 6 × 10^22 atoms contributes 6 × 10^22 electrons.